# Commuting integrable maps from a deformed D4 cluster algebra

### v1

@@ 2 Deformation of the $\mathrm{D}_{4}$ cluster map
where $c_{j}$ are parameters to be determined. Then imposing the requirement that $\hat{\varphi}^{*}(\tilde{K})=\tilde{K}$, leads us to necessary and sufficient conditions for the map $\hat{\varphi}$ to be integrable with an invariant of this form, as follows.
@@ 2.2 Rational elliptic surface
where the third factor $\ell_{1}$ is linear in $w$; so each factor defines a rational curve. It turns out that the two functions above become equivalent to one another when $\mathrm{b}_{3}=1$, upon identifying the remaining parameters $\mathrm{b}_{1}$ and $\mathrm{b}_{4}$ in each case, although the associated maps $\hat{\varphi}_{1},\hat{\varphi}_{2}$ remain distinct from one another. Similarly, the second factor in the denominator of the j-invariant gives a double pole at ${\kappa}=-{\alpha}{\beta}-3$, and for this value of $\kappa$} the equation of the curve decomposes as
@@ 3.2 The commuting map $\hat{\chi}$
    Hence we have a well defined map of the plane, and the fact that $\hat{\chi}$ is birational essentially follows from the birationality of the cluster map $\chi$, which is composed of mutations (which are involutions) together with a permutation. Now we can also check directly that $\hat{\varphi}\cdot\hat{\chi}=\hat{\chi}\cdot\hat{\varphi}$, so the maps commute, and by lifting this implies that the cluster maps $\psi$ and $\chi$ also commute with one another, i.e.
@@ 3.3 Tau functions on the ${\mathbb{Z}}^{2}$ lattice
$\tilde{{\bf x}}_{m,n}=\chi^{m}\psi^{n}(\tilde{{\bf x}})=(T_{m,n-2},T_{m-1,n},X_{m,n},T_{m+1,n+2},T_{m,n},T_{m,n+1},T_{m,n+2},T_{m,n-1}).$
Note that, due to the different weights of the equations (3.8), we are required to use two different tau functions on the lattice: $T_{m,n}$, which has weight 1, and $X_{m,n}$, which has weight 2. By considering the action of the symplectic maps $\hat{\chi}$, $\hat{\varphi}$ on the plane with coordinates ${\bf u}=(z,w)$, we can also define a corresponding sequence of combined iterates of these two maps, denoted by
@@
$\begin{array}[]{rclcl}T_{m,n+3}T_{m,n-2}&=&{\alpha}{\beta}\,T_{m,n+1}T_{m,n}&+&X_{m,n},\\ X_{m,n+1}X_{m,n}&=&T_{m+1,n+2}T_{m,n+3}T_{m,n-1}T_{m-1,n}&+&{\alpha}{\beta}\,T_{m,n+2}T_{m,n+1}^{2}T_{m,n},\\ T_{m+1,n+2}T_{m-1,n+1}&=&{\alpha}\,T_{m,n+2}T_{m,n+1}&+&X_{m,n+1},\\ T_{m+1,n+3}T_{m-1,n}&=&{\beta}\,T_{m,n+2}T_{m,n+1}&+&X_{m,n+1},\\ T_{m+1,n}X_{m,n}&=&{\alpha}\,T_{m+1,n+2}T_{m,n-1}T_{m,n}&+&T_{m+1,n+1}T_{m,n+2}T_{m,n-2},\\ T_{m+1,n-1}T_{m,n+2}&=&{\alpha}\,T_{m,n-1}T_{m+1,n+2}&+&T_{m,n+1}T_{m+1,n},\\ X_{m+1,n}T_{m,n-2}&=&{\beta}\,T_{m+1,n}T_{m+1,n-1}T_{m,n}&+&T_{m+1,n+2}T_{m+1,n-2}T_{m,n-1}.\end{array}$ (3.15)
@@ 3.4 Degree growth of cluster maps
The Laurent property means that each of the tau functions $T_{m,n},X_{m,n}$ satisfying the lattice system (3.15) is a Laurent polynomial in the initial seed variables (3.13) with coefficients in ${\mathbb{Z}}[{\alpha},{\beta}]$. Hence we can write each of these cluster variables in the form
@@
where $\mathrm{N}_{m,n},\mathrm{N}^{*}_{m,n}\in{\mathbb{Z}}[\hat{{\bf x}}$] are polynomials in the extended cluster variables $\hat{{\bf x}}$ that are not divisible by any $\tilde{x}_{j}$, $1,\leq j\leq 8$, and the denominators are monomials in the seed variables $\tilde{{\bf x}}$, with the exponent of each $\tilde{x}_{j}$ being specified by the $j$th component of the d-vectors ${\bf d}_{m,n},{\bf e}_{m,n}\in{\mathbb{Z}}^{8}$, respectively. Due to the homogeneity properties of these tau functions, as observed above, the degrees of each numerator and denominator are completely determined by the d-vectors. Then, in order to calculate the growth of degrees, we can use the basic fact, pointed out in [4], that the d-vectors for cluster variables satisfy the tropical (max-plus) version of the exchange relations, in which the frozen variables do not make any contribution. This immediately yields the following result, by taking the appropriate max-plus analogues of (3.15).
@@
  • Proof:
    The $n$ part of the statement, namely the fact that the d-vectors lie in the kernel of $\tilde{\cal L}$, was already given in Theorem 5.11 of [7], together with the coefficient vector $\bf a$} multiplying the leading $n^{2}$ growth in this lattice direction. The proof of the rest of the statement follows a very similar pattern, and also closely resembles the proofs of analogous results on the solution of a tropical system associated with a deformed $\mathrm{A}_{3}$ map (cf. Lemma 2.10 and Theorem 2.11 in [7]). The main point is that the solution of (3.17) on ${\mathbb{Z}}^{2}$ is completely determined by fixing a tropical seed, with the specific choice (3.18) being the one that corresponds to (3.13). Then the expression (3.22) can be rewritten as
@@
    and then note that by the 4-periodicity in Lemma 3.6 we have
@@ 4.1 Space of initial conditions
Table 1: Curves on the surface $\cal X$}. The numbers in the definition correspond to those in Figures 2 and 3.
@@
for a basis of $\operatorname{Pic}{\cal X}$, where [][-] denotes a class in $\operatorname{Pic}{\cal X}$. The choice is made because the motion of $E_{j}$ and $D_{j}$ is cyclic, and the motion of $C_{1}$ and $C_{7}$ is complicated. The intersection matrix on $\operatorname{Pic}{\cal X}$ for this basis is
@@ 4.2 Degree growth
Firstly, we calculate the degree growth and the leading coefficient for the deformed D4 map, i.e., $\hat{\varphi}$. Calculating the motion of each curve, we have the action of $\hat{\varphi}$ on $\operatorname{Pic}{\cal X}$ as follows:
@@
However, because of the large size of $M^{n}_{\hat{\varphi}}$, it is not easy to directly calculate the Jordan basis for $M^{n}_{\hat{\varphi}}$ by software. Here, we will use the general theory from [10].
@@
acts on $\mathbb{C}^{14}$ as a projection to the generalized eigenspace for $1$. A direct calculation shows that the $9$-th column of the matrix
@@
is non-zero. Let $v_{3}\in\mathbb{Z}^{14}$be the $9$-th column of $(M_{\hat{\varphi}}+I_{14})(M^{2}_{\hat{\varphi}}+I_{14})(M^{2}_{\hat{\varphi}}+M_{\hat{\varphi}}+I_{14})$, i.e.,
@@
Secondly, we calculate the degree growth and the leading coefficient for $\hat{\chi}$ in the same way. The action of $\hat{\chi}$ on $\operatorname{Pic}{\cal X}$ is
@@
Thirdly, we calculate the degree growth and the leading coefficient for the composition $\hat{\varphi}\cdot\hat{\chi}$. Let
@@
A direct calculation shows that the $9$-th column of the matrix
@@
Let us show that the maps $\hat{\varphi}$ and $\hat{\chi}$ are independent in the same way as in [7, Appendix A]. Define the canonical height $\hat{h}$ and associated bilinear pairing <,><-,-> on the Mordell-Weil group in the usual way, and let $P_{1}$ and $P_{2}$ be the translations on $\cal X$} (considered as an elliptic curve over $\mathbb{C}({\kappa})$) corresponding to $\hat{\varphi}$ and $\hat{\chi}$, respectively. Then we have
@@
Since this matrix is non-singular, the maps $\hat{\varphi}$ and $\hat{\chi}$ are independent.
@@ 4.3 QRT map
Then we can defined the composition
@@
Surprisingly, if rewritten back in the chart $(z,w)$, the QRT map $\varphi_{\mathrm{QRT}}$ coincides with the deformed D4 map $\hat{\varphi}$.
@@ Acknowledgments
for details. ANWH is grateful to BIMSA for supporting his stay there during the RTISART conference in July 2024, and to the Graduate School of Mathematical Sciences, University of Tokyo for the invitation to speak at the FoPM Symposium in February 2025, where this paper was first conceived. He would also like to thank Tetsushi Ito (Kyoto) for supporting his subsequent visit to Japan in June, where the Number Theory & Integrable Systems workshop in Kobe provided a further opportunity for us to work on the paper together.

### v2

@@ 2 Deformation of the $\mathrm{D}_{4}$ cluster map
where $c_{j}$ are parameters to be determined. Then imposing the requirement that $\hat{\varphi}^{*}(\tilde{K})=\tilde{K}$ leads us to necessary and sufficient conditions for the map $\hat{\varphi}$ to be integrable with an invariant of this form, as follows.
@@ 2.2 Rational elliptic surface
where the third factor $\ell_{1}$ is linear in $w$; so each factor above defines a rational curve. Similarly, the second factor in the denominator of the j-invariant gives a double pole at ${\kappa}=-{\alpha}{\beta}-3$, and for this value of $\kappa$} the equation of the curve decomposes as
@@ 3.2 The commuting map $\hat{\chi}$
Hence we have a well defined map of the plane, and the fact that $\hat{\chi}$ is birational essentially follows from the birationality of the cluster map $\chi$, which is composed of mutations (which are involutions) together with a permutation. Now we can also check directly that $\hat{\varphi}\cdot\hat{\chi}=\hat{\chi}\cdot\hat{\varphi}$, so the maps commute, and by lifting we find that the cluster maps $\psi$ and $\chi$ also commute with one another, i.e.
@@ 3.3 Tau functions on the ${\mathbb{Z}}^{2}$ lattice
$\tilde{{\bf x}}_{m,n}=\chi^{m}\psi^{n}(\tilde{{\bf x}})=(T_{m,n-2},T_{m-1,n},X_{m,n},T_{m+1,n+2},T_{m,n},T_{m,n+1},T_{m,n+2},T_{m,n-1}).$ (3.14)
The initial cluster (3.13) corresponds to the stencil of lattice points shown in Figure 2, where the 7 crosses represent the tau functions $T_{ij}$, and the box represents $X_{0,0}$, placed at the origin in ${\mathbb{Z}}^{2}$; and each cluster (3.14) corresponds to the same stencil but shifted $m$ steps to the right/left and $n$ steps up/down, according to whether $m,n$ are positive/negative, respectively. Note that, due to the different weights of the equations (3.8), we are required to use two different tau functions on the lattice: $T_{m,n}$, which has weight 1, and $X_{m,n}$, which has weight 2. By considering the action of the symplectic maps $\hat{\chi}$, $\hat{\varphi}$ on the plane with coordinates ${\bf u}=(z,w)$, we can also define a corresponding sequence of combined iterates of these two maps, denoted by
@@
$T_{m,n+3}T_{m,n-2}$ $={\alpha}{\beta}\,T_{m,n+1}T_{m,n}$ $+\quad$ $X_{m,n},$ ()
$X_{m,n+1}X_{m,n}$ $=T_{m+1,n+2}T_{m,n+3}T_{m,n-1}T_{m-1,n}$ $+\quad$ ${\alpha}{\beta}\,T_{m,n+2}T_{m,n+1}^{2}T_{m,n},$ ()
$T_{m+1,n+2}T_{m-1,n+1}$ $={\alpha}\,T_{m,n+2}T_{m,n+1}$ $+\quad$ $X_{m,n+1},$ ()
$T_{m+1,n+3}T_{m-1,n}$ $={\beta}\,T_{m,n+2}T_{m,n+1}$ $+\quad$ $X_{m,n+1},$ ()
$T_{m+1,n}X_{m,n}$ $={\alpha}\,T_{m+1,n+2}T_{m,n-1}T_{m,n}$ $+\quad$ $T_{m+1,n+1}T_{m,n+2}T_{m,n-2},$ ()
$T_{m+1,n-1}T_{m,n+2}$ $={\alpha}\,T_{m,n-1}T_{m+1,n+2}$ $+\quad$ $T_{m,n+1}T_{m+1,n},$ ()
$X_{m+1,n}T_{m,n-2}$ $={\beta}\,T_{m+1,n}T_{m+1,n-1}T_{m,n}$ $+\quad$ $T_{m+1,n+2}T_{m+1,n-2}T_{m,n-1}.$ ()
@@
At any fixed lattice point $(m,n)\in{\mathbb{Z}}^{2}$, by starting from a seed cluster $\tilde{{\bf x}}_{m,n}$ given by (3.14), corresponding to the points in the stencil shown in Figure 2, we can identify the cluster variables via
$\tilde{{\bf x}}_{m,n}=(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3},\tilde{x}_{4},\tilde{x}_{5},\tilde{x}_{6},\tilde{x}_{7},\tilde{x}_{8}),$
just as in (3.13) for the case $m=0=n$, and further extend this cluster by adjoining the frozen variables $\tilde{x}_{9}={\alpha},\tilde{x}_{10}={\beta}$. To compute the action of $\chi$ on this cluster, we begin by applying the mutation $\tilde{\mu}_{2}$, which corresponds to the equation (()3.4) after shifting down the index $n\to n-1$, and produces the new cluster variable $\tilde{x}_{2}^{\dagger}=T_{m+1,n+1}$. Subsequently, we apply (in order) the mutations $\tilde{\mu}_{3},\tilde{\mu}_{7},\tilde{\mu}_{6},\tilde{\mu}_{1},\tilde{\mu}_{8}$, corresponding to the equations (()3.4), (()3.4), (()3.4) again but with $n\to n-1$, (()3.4), (()3.4) but with $m\to m+1$ and $n\to n-1$, which produce the new cluster variables $\tilde{x}_{3}^{\dagger}=T_{m+1,n}$, $\tilde{x}_{7}^{\dagger}=T_{m+1,n-1}$, $\tilde{x}_{6}^{\dagger}=T_{m+1,n-2}$, $\tilde{x}_{1}^{\dagger}=X_{m+1,n}$, $\tilde{x}_{8}^{\dagger}=T_{m+2,n+2}$, respectively. Finally, we apply the permutation $(\rho^{\dagger})^{-1}$, which from (3.7) yields
$\chi(\tilde{{\bf x}}_{m,n})=(T_{m+1,n-2},T_{m,n},X_{m+1,n},T_{m+2,n+2},T_{m+1,n},T_{m+1,n+1},T_{m+1,n+2},T_{m+1,n-1})=\tilde{{\bf x}}_{m+1,n},$
that is, the new seed with $m\to m+1$, which corresponds to shifting all the crosses and the box that make up the stencil in Figure 2 one place to the right. Moreover, by the argument used to prove Theorem 3.2, the pair of quantities $(z_{m+1,n},w_{m+1,n})$ given by the appropriate ratios of these cluster variables, as defined by the expressions for $z_{m,n}$ and $w_{m,n}$ in (3.15) with $m\to m+1$, are precisely the image of $(z_{m,n},w_{m,n})$ under the map $\hat{\chi}$, that is
$\hat{\chi}\Big((z_{m,n},w_{m,n})\Big)=(z_{m+1,n},w_{m+1,n}).$
Similarly, starting from the same seed cluster $\tilde{{\bf x}}_{m,n}$ with the two frozen variables ${\alpha},{\beta}$, we can compute the action of the cluster map $\psi$, as in (3.5). Then we see that taking the mutations $\tilde{\mu}_{1},\tilde{\mu}_{3},\tilde{\mu}_{4},\tilde{\mu}_{2}$, in that order, corresponds to the equations (()3.4), (()3.4), (()3.4), (()3.4), which produce the new cluster variables $\tilde{x}_{1}^{\prime}=T_{m,n+3}$, $\tilde{x}_{3}^{\prime}=X_{m,n+1}$, $\tilde{x}_{4}^{\prime}=T_{m-1,n+1}$, $\tilde{x}_{2}^{\prime}=T_{m+1,n+3}$, respectively. Thus, after applying the permutation $\rho^{-1}$, we obtain the new cluster, namely
$\psi(\tilde{{\bf x}}_{m,n})=(T_{m,n-1},T_{m-1,n+1},X_{m,n+1},T_{m+1,n+3},T_{m,n+1},T_{m,n+2},T_{m,n+3},T_{m,n})=\tilde{{\bf x}}_{m,n+1},$
which is the new seed with $n\to n+1$, and this corresponds to moving all the crosses and the box in Figure 2 one place up. The result of Theorem 5.6 in [7] also shows that suitable ratios of the tau functions from the above cluster produce the pair $(z_{m,n+1},w_{m,n+1})$, given by applying the shift $n\to n+1$ to the formulae for $z_{m,n}$ and $w_{m,n}$ in (3.15), and this coincides with the image of $(z_{m,n},w_{m,n})$ under the symplectic map $\hat{\varphi}$, that is
$\hat{\varphi}\Big((z_{m,n},w_{m,n})\Big)=(z_{m,n+1},w_{m,n+1}).$
To complete the proof, it is enough to note the fact that, from the commutativity of the two cluster maps, as in (3.12), it is consistent to extend the system (()3.4)-(()3.4) to all $(m,n)\in{\mathbb{Z}}^{2}$, and then from (3.15) the solution of this lattice system provides a simultaneous solution for the pair of iterated commuting maps $\hat{\chi}$ and $\hat{\varphi}$. ∎
@@ 3.4 Degree growth of cluster maps
The Laurent property means that each of the tau functions $T_{m,n},X_{m,n}$ satisfying the lattice system (()3.4)-(()3.4) is a Laurent polynomial in the initial seed variables (3.13) with coefficients in ${\mathbb{Z}}[{\alpha},{\beta}]$. Hence we can write each of these cluster variables in the form
@@
where $\mathrm{N}_{m,n},\mathrm{N}^{*}_{m,n}\in{\mathbb{Z}}[\hat{{\bf x}}$] are polynomials in the extended cluster variables $\hat{{\bf x}}$ that are not divisible by any $\tilde{x}_{j}$, $1,\leq j\leq 8$, and the denominators are monomials in the seed variables $\tilde{{\bf x}}$, with the exponent of each $\tilde{x}_{j}$ being specified by the $j$th component of the d-vectors ${\bf d}_{m,n},{\bf e}_{m,n}\in{\mathbb{Z}}^{8}$, respectively. Due to the homogeneity properties of these tau functions, as observed above, the degrees of each numerator and denominator are completely determined by the d-vectors. Then, in order to calculate the growth of degrees, we can use the basic fact, pointed out in [4], that the d-vectors for cluster variables satisfy the tropical (max-plus) version of the exchange relations, in which the frozen variables do not make any contribution. This immediately yields the following result, by taking the appropriate max-plus analogues of (()3.4)-(()3.4).
@@
The $n$ part of the statement, namely the fact that the d-vectors lie in the kernel of $\tilde{\cal L}$, was already given in Theorem 5.11 of [7], together with the coefficient vector $\bf a$} multiplying the leading $n^{2}$ growth in this lattice direction. The proof of the rest of the statement follows a very similar pattern, and also closely resembles the proofs of analogous results on the solution of a tropical system associated with a deformed $\mathrm{A}_{3}$ map (cf. Lemma 2.10 and Theorem 2.11 in [7]). The main point is that the solution of (3.18) on ${\mathbb{Z}}^{2}$ is completely determined by fixing a tropical seed, with the specific choice (3.19) being the one that corresponds to (3.13). Then we see that, on the one hand, the expression (3.23) can be rewritten as
@@
while on the other hand, by the 4-periodicity in Lemma 3.6, we have
@@ 4.1 Space of initial conditions
Table 1: Curves on the surface $\cal X$}. The numbers in the definition correspond to those in Figures 3 and 4.
@@
for a basis of $\operatorname{Pic}{\cal X}$, where [][-] denotes a class in $\operatorname{Pic}{\cal X}$. The choice is made because the motion of $E_{j}$ and $D_{j}$ is cyclic, and the motion of $C_{1}$ and $C_{7}$ is complicated. The intersection matrix on $\operatorname{Pic}{\cal X}$ in this basis is
@@ 4.2 Degree growth
Proposition 4.1.
The degree growth of the map $\hat{\varphi}$ is quadratic, and its leading coefficient is $\frac{2}{3}$.
Proof.
Calculating the motion of each curve, we have the action of $\hat{\varphi}$ on $\operatorname{Pic}{\cal X}$ as follows:
@@
However, because of the large size of $M^{n}_{\hat{\varphi}}$, it is not easy to directly calculate the Jordan basis for $M^{n}_{\hat{\varphi}}$ with computer algebra. Here, we will use the general theory from [10].
@@
acts on $\mathbb{C}^{14}$ as a projection to the generalized eigenspace for $1$. A direct calculation shows that the $9$th column of the matrix
@@
is non-zero. Let $v_{3}\in\mathbb{Z}^{14}$be the $9$th column of $(M_{\hat{\varphi}}+I_{14})(M^{2}_{\hat{\varphi}}+I_{14})(M^{2}_{\hat{\varphi}}+M_{\hat{\varphi}}+I_{14})$, i.e.,
@@
Proposition 4.2.
The degree growth of the map $\hat{\chi}$ is quadratic, and its leading coefficient is $\frac{14}{3}$.
Proof.
The action of $\hat{\chi}$ on $\operatorname{Pic}{\cal X}$ is
@@
Proposition 4.3.
The degree growth of the composition map $\hat{\varphi}\cdot\hat{\chi}$ is quadratic, and its leading coefficient is $4$.
Proof.
Let
@@
A direct calculation shows that the $9$th column of the matrix
@@
Theorem 4.4.
The maps $\hat{\varphi}$ and $\hat{\chi}$ are independent.
Proof.
The proof follows the same argument as in [7, Appendix A]. Define the canonical height $\hat{h}$ and associated bilinear pairing <,><-,-> on the Mordell-Weil group in the usual way, and let $P_{1}$ and $P_{2}$ be the translations on $\cal X$} (considered as an elliptic curve over $\mathbb{C}({\kappa})$) corresponding to $\hat{\varphi}$ and $\hat{\chi}$, respectively. Then we have
@@
Since this matrix is non-singular, the maps $\hat{\varphi}$ and $\hat{\chi}$ are independent. ∎
@@ 4.3 QRT map
Then we can define the composition
@@
The following coincidence is remarkable.
Proposition 4.6.
If rewritten back in the chart $(z,w)$, the QRT map $\varphi_{\mathrm{QRT}}$ coincides with the deformed D4 map $\hat{\varphi}$.
Proof.
By direct calculation. ∎
@@ Acknowledgments
for details. ANWH is grateful to BIMSA for supporting his stay there during the RTISART conference in July 2024, and to the Graduate School of Mathematical Sciences, University of Tokyo for the invitation to speak at the FoPM Symposium in February 2025, where this paper was first conceived. He would also like to thank Tetsushi Ito (Kyoto) for supporting his subsequent visit to Japan in June, where the Number Theory & Integrable Systems workshop in Kobe provided a further opportunity for us to work on the paper together. The authors would like to thank an anonymous referee for helpful suggestions which improved the clarity of the paper.